From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: August 6, 2009

Participants: Daniel, Brian, Bill, Beni, Simon, Rich, Fernando, Mark I., Mark S., Eugene, Roger, David, Micah, Chuck

Next Meeting: August 13, 2009 @ 1:30 p.m.

General Announcements

 - Today was Rich Rines last day with us. The Hall D gang wants to thank
   him for his hard work on the FDC project this summer and to wish him
   well on his studies at UMass.

General Construction Updates

 - The parts for the assembly jig on which the detector stack will be
   constructed will be put together as there is time. This should probably
   happen in the next week as it has been moved to our short term work
 - Bill has begun work on the FDC stack assembly document. The draft will
   be made available after he completes a bit more work on it.  The document
   will be fleshed out during the assembly process with relevant photographs.
 - The composite frames for the two gas windows have been completed. Sealing
   of the I.D. and O.D. of the frames has been completed, with the sealing
   of the bolt holes remaining. The tensioning of the gas windows and the 
   lamination to the composite rings will be done as we have time in the next 
   couple of weeks. 
 - Mark S. has begun construction of the ground cable pig tails that will
   be used to attach to the cathode planes. He will make up about 20 of
   the cables initially and make up more as they are needed for the 
   full-scale prototype construction.
 - A final decision has been made on the conducting epoxy that we will use
   in the prototype. Mark S. has prepared the order.
 - We still need to follow up with Steve Christo to find out the status
   of the wire order and the status of the qualification of the vendors.
 - More space has been taken in EEL 126 after one of the large dark boxes
   was moved away. We have set up a humidity-controlled tent in the space
   as the humidity in the room was measured at about 60-65% (even though
   our cheap-o Radio Shack devices were reading 40-45%.

Wire Frame Update

 - Rich has done what he could to remove the solder from frames P4 and P3
   and is currently working on cleaning frame P2. This work should be 
   completed by the end of the day today. He has also been checking over
   all components on the boards to flag those that are missing, are
   problematic, or have the wrong value. Armen Stepanyan from the Fast
   Electronics Group has been coming in and doing the repair work as needed.
 - Rich has found several HV busses on each wire frame (P4 and P3) that seem 
   to persistently draw about 1 uA of current. Whether this is related to the 
   relative humidity in EEL 126 is unclear. The boards have been cleaned
   extensively and checked under the microscope for other problems. We 
   decided to dry the frame out in the humidity-controlled tent over the
   weekend and to retest next week.  
 - The last wire plane that needs checking and repairs is P1. This work will
   not happen in the short term as only 3 frames are needed for the
   full-scale prototype and we don't have the manpower to work on this now.
 - Rich will provide a write-up on his QA check procedures and make this
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We will include
   the information on the wire electroplating process shortly.
 - The parts of the jig to use for the wire electroplating operation
   are in our hands. They will be assembled as we find time.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   when he gets a chance.

Cathode Update

 - Progress updates:
     * One of our completed cathode boards is really contaminated with
       epoxy near the two seams. The epoxy manufacturer sent us a solvent 
       that they believe should allow the epoxy to be removed. Bill did a
       test that effectively removed the epoxy by also discolored the
       Cu surface on his test sample. This was due to the fact that the
       solvent was not effectively neutralized. Bill will talk to the
       manufacturer to find out what they recommend. It may be that we
       can save this cathode after all.
     * It was found that the cathodes were not sitting flat on the flatness
       system. Bill came up with a work around that fixed the problem
       wonderfully. Flatness measurements can now proceed.
     * Ground planes have been attached to the two transfer rings and one
       has been mounted on the tensioner. We need to tension this one and
       use our experience to define the nominal tension. Lamination to the
       cathode should proceed following by glueing down of the ground
       flaps and attachment of the ground pigtail.
 - Micah found a problem with the speed settings and laser head settings
   on the flatness system and has made fixes. DSC wants to see the nominal
   set of plots made for a given scan and needs to devise a criterion for
   cathode flatness acceptance.
 - Bill is devising a work plan for the prototype construction that he
   will give to Chuck to put into a schedule for tracking.
 - The Fast Electronics group completed the soldering and initial QA checks
   on the rigid-flex assemblies. Fernando stated that the QA checks done
   were visual only. An electrical QA check is essential and he will come
   up with a test station and test plan that we can use within the next
   couple of weeks. 
 - Mark S. will begin test soldering of the rigid-flex assemblies to
   test cathode pieces to finalize the soldering procedures. He will work
   under the direction of Roger and Fernando.
 - Bill has been working on the cathode construction document building in
   the pictures that are being taken. The current draft of the document 
   is located at /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We need
   feedback from folks involved in the process to be sure that all steps
   are included and accurately detailed.
 - We are standing by waiting for information on the PR for the 2-um board
   material. We have learned that the order has gone out, but the Japanese
   company has still not indicated whether they will fill the order. A
   second company that can produce the board material has been identified.
   Roger has been in contact with the U.S. representative of this company
   and believes that we will have an answer by the end of this month.
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. Progress is slow and steady. Safety paperwork is now
   in place and gas is flowing to all chambers. We have seen a preliminary
   FADC spectrum from the cathodes and TDC spectrum from the wires on the
   small-scale prototype. Beni will work to reproduce the "picket fence"
   plot that Simon has produced to quantify the resolution last year.
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. We decided to have a separate meeting to discuss what work
   we would like to do on the small-scale prototype.
 - Beni has started a bench test plan for the full-scale prototype and
   is presently working on it after incorporating feedback. The current
   information is contained on the FDC wiki page under the heading 
   "FDC test cosmic ray test stand".
 - The parts for the support/rotation of the full-scale prototype in the 
   cosmic ray stand will be assembled as we have time.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.